# Role of Interlayer Shear Phonons on the Lattice Symmetry Switch in a Transition-Metal Dichalcogenide

### v1

@@ Abstract
Coherent phonon control using ultrashort pulse trains is the key to realizing structural phase transitions in solids by non-thermal pathways. By combining double-pulse excitation and time-resolved second harmonic generation techniques under high-density electronic excitation in a 2D layered material, WTe2, we demonstrate that the lattice symmetry switching from the Weyl semimetallic Td to the semimetallic 1T′ phases is independent of the amplitude of the coherent interlayer shear phonons. This finding provides new insights into the mechanisms for symmetry switching that electronic excitation-driven shear sliding plays a dominant role.
@@
Recently, a feasible system for optical control of coherent PIPT involves ultrafast lattice symmetry switching between the semimetallic 1T′ (centrosymmetric) and Weyl semimetallic Td (non-centrosymmetric) phases of layered transition-metal dichalcogenides (TMDs) such as and [21, 22, 23, 24]. Notably, driving a coherent interlayer shear phonon displacement provides a means to control the lattice symmetry and allow periodic alternation of adjacent structural phases at sub-THz frequencies. To realize coherent PIPT for optical control, it is essential to elucidate the dynamics of incoherent and coherent PIPT in TMDs, focusing on the effect of interlayer shear displacement. Additionally, under high-density electronic excitation conditions exceeding several by a single pulse, the intended phase transition may be suppressed by sample damage or saturable absorption [25, 26, 27]. These issues can arise not only in controlling lattice symmetry but also in realizing optical control of other coherent PIPTs, such as spin states [28], and must be addressed.
In this paper, we address these issues through experimental studies on lattice symmetry change under high-density electronic excitation using time-resolved second harmonic generation (TR-SHG), a sensitive probe for tracking the symmetry change [21, 22, 29, 30, 31]. By combining double-pulse excitation and TR-SHG methods in a two-dimensional layered material system, WTe2, we show that the symmetric switch from the broken-inversion Weyl semimetal Td phase to inversion-symmetric semimetal 1T′ phase is independent of the amplitude of the coherent shear phonons, which has been believed to play a central role.
The sample used in this study was a bulk single crystal of (001) surface Td-WTe2 (from HQ Graphene) with a thickness of 100 m. The sample surface was cleaved with an adhesive tape prior to measurements. SHG arises from a non-zero second-order susceptibility in materials without inversion symmetry [32, 33, 34]. In WTe2, the Td phase has no inversion symmetry, whereas the 1T′ phase does. Consequently, the Td phase exhibits SHG, while the 1T′ does not [21].
@@
Before TR-SHG measurements, static SHG measurements were performed to clarify the SHG response from Td-WTe2. Figure 1(b) shows the spectrum observed for the incoming probe photon (1230 nm) and the resulting SHG signal (615 nm) emitted from the sample. Figure 1(c) shows the dependence of the polarization angle on the 1230 nm light pulses measured without pump light to investigate the response of SHG to crystal orientation. SHG intensity reaches its maximum for the -axis polarization and minimum for the -axis polarization. For TR-SHG measurements, the polarization was set to the -axis to maximize the SHG intensity. Here, the intense probe (3.2 ) was used to obtain a sufficient signal-to-noise ratio due to the weak second harmonic effect. It is noted that the intense probe does not influence PIPT because the fluence was selected within the region where the intensity of SHG was proportional to the square of the intensity of the incident probe , i.e., up to 4 mJ/cm2 (see details in Supplemental Material (SM))[35], as expected for a second-order nonlinear process. If the probe contributes to the PIPT, the SHG should decrease following the lattice symmetry changes [36, 22]. However, this is not the case here. Furthermore, according to the absorption coefficient obtained from our spectroscopic ellipsometry measurement, the effective photogenerated carrier density can be estimated to be 4.3 times larger for the pump than for the probe at the same fluence level (see SM) [35]. This additional measurement further supports the conclusion that the intense probe does not influence PIPT.
@@
To explore the contribution of electronic excitation and shear phonon oscillation in structural change, we performed coherent control of the interlayer shear phonon. Figure 2 shows the time-domain signal of the SHG intensity change () obtained under double-pulse excitation together with the case of single-pulse excitation at 1.5 mJ/cm2 (see the inset). As seen in the inset, the SHG intensity decreases just after photoexcitation (2 ps) and then exhibits a long-lived coherent oscillation for tens of picoseconds. The frequency of the long-lived coherent oscillation was found to be 0.24 THz, matching the interlayer shear phonon assigned as the mode [21, 37, 29]. The time interval between the first and second pulses is 1.5 (= 6.25 ps) and 2 (= 8.33 ps) of the interlayer shear phonon at the identical total pump fluence of 4.6 mJ/cm2, where = 4.17 ps is the period of the interlayer shear phonon.
@@
The decrease in SHG intensity indicates that the lattice symmetry changes from the Td toward 1T′ phases [21]. When the value of is an integral multiple of the phonon period, the oscillation amplitude of the coherent phonon generated by the first pulse is enhanced by the second pulse [11]. In contrast, when it is a half-integer multiple of the phonon period, it is canceled by the second pulse. As expected, the shear phonon was canceled when =1.5 and enhanced when =2. From Fig. 2, the only difference between the two cases is whether or not the oscillation of shear phonon, and no significant difference in the magnitude of was observed after 10 ps. Therefore, in the symmetry change from the Td toward 1T′ phases under the near-infrared (NIR) photon pump (800 nm), the effect of electronic excitation is a more dominant factor than the displacement of the shear phonon. In other words, the excitation of the coherent shear phonon can be independent of the formation of the metastable phase. Although decoupling of the shear phonon from the symmetry change from the Td toward 1T′ phases was suggested by suppressing the shear phonon with lattice defects [37], our results indicate more direct evidence by using coherent control of the phonon amplitude.
@@
To gain more insight into the effects of coherent control of the phonon amplitude on the possible lattice symmetry change from the Td toward 1T′ phases, we present time-domain SHG signals for both constructive ( = 8.33 ps) and destructive ( = 6.25 ps) excitation at the total pump fluence from 1.6 (1.8) mJ/cm2 to 10.6 mJ/cm2 in Fig. 3. The amplitude of the shear phonon is enhanced by constructive excitation, as shown in Fig. 3(a), whereas it is canceled by destructive excitation as shown in Fig. 3(b). In both cases, the SHG intensity decreases as the total pump fluence increases.
@@
Figure 4 shows the pump fluence dependence of obtained from the fitting in Fig. 3. As seen in Fig. 4, higher pump fluence leads to a more significant decrease in the SHG intensity for both single- and double-pulse excitation. Moreover, even in the fluence dependence, the SHG intensity is significantly lower for double-pulse excitation than for single-pulse excitation. This suggests that saturable absorption was suppressed by multiple-pulse excitation [38, 25, 15] and/or that the structure is more likely to change if a second pump pulse is applied after the first has already modified the structure [20]. Although there has been a Landau theory to account for the temperature dependence of the SHG intensity [39], to our knowledge, an appropriate model for the fluence dependence will currently not be available under double-pulse excitation. Thus, the fluence dependence can be adapted to two linear functions with different slopes, implying a change of the potential energy surface [40, 1], e.g., a shift of the potential energy minimum toward the 1T′ phase, under PIPT. In this case, the critical fluence will be higher for double-pulse excitation (6 mJ/cm2) than for single-pulse excitation (4 mJ/cm2). More experimental and theoretical studies will be required to fully understand the fluence dependence found in Fig. 4.
@@
To evaluate the impact of thermal effects during the phase transition from the Weyl semimetal Td to semimetal 1T′ phases, we have calculated the electron and lattice temperatures using the two-temperature model (TTM) [41] under double-pulse excitation as well as incorporating possible cumulative effects [42]. TTM, which applies to zero-gap metallic systems, can be used to examine the effect of elevated lattice temperatures. The TTM results for the two typical pump fluences in our experiment (7.0 mJ/cm2 and 9.2 mJ/cm2) are shown in SM[35]. Although the TTM analysis can only be used as a crude estimate, after the arrival of the second pulse, the lattice temperature () exceeds the transition temperature ( = 565 K[43] or 613 K [44]) at the time delay of 10 ps for 9.2 mJ/cm2. On the other hand, is below for 7.0 mJ/cm2. Interestingly, this critical fluence (7.0 mJ/cm2) is close to that observed for the depletion of the SHG intensity in Fig. 4, implying that the flatter part for 6 mJ/cm2 would need to be further addressed focusing on, in particular, early-time dynamics (first several picoseconds) where the shear phonon displacements facilitate the symmetry switch. Note that the actual temperature increase estimated may slightly differ from the TTM results because the TTM maintains homogeneous absorption, no multi-photon absorption, both of which could occur under high-density photoexcitation, this is the case here. Nevertheless, we can conclude that the lattice symmetry switching from the Weyl semimetallic Td to the semimetallic 1T′ phases is not promoted by the lattice temperature rise at the total pulse fluence of 7 mJ/cm2.
@@
We argue that the lattice symmetry switch from the Td to the 1T′ phase is driven either by interband or intraband electronic excitation. The former promotes electrons from bonding to anti-bonding states, increasing the charge carrier density . In contrast, the latter intraband transition results in an increase in the electron temperature (or electron acceleration) without changing and can preferably be achieved by irradiating far-infrared or THz pulses [21]. Both excitation processes can promote carrier (hole) doping and induce shear sliding of atomic layers toward the 1T′ phase [9]. Under NIR (800 nm) conditions, the photoexcited carriers in WTe2 are generated by interband transitions along and directions in the electronic band structure [45]. In fact, we measured the imaginary part () of the complex refractive index of WTe2 by spectroscopic ellipsometry, and found that the 800 nm ( 1.55 eV) pump light used in the present study was resonant with 1.56 eV (see details in SM)[35], which can generate the photoexcited carrier via the interband transition. Thus, the increase in charge carrier density, e.g., 1.26 1021/cm3 (at 1.6 mJ/cm2) due to interband transitions plays a dominant role in the symmetry switch.
@@
This study demonstrates that coherent control of the shear phonon amplitude can suppress its effect, isolating the impact of electronic excitation. We find that inversion-symmetry changes in the crystal structure mainly result from electronic excitation, i.e., incoherent PIPT [1]. Although previous attempts have used coherent shear phonons to induce structural transitions, we successfully decouple shear phonon motion from the Td-to-1T′ phase switch in WTe2 through destructive excitation. Since the TR-SHG technique is not a method that directly measures lattice displacement, in particular, atomic motions during this early time window (several picoseconds), it will be important to examine the present results using direct methods such as ultrafast electron diffraction [17, 21] and time-resolved x-ray diffraction [46, 47], which can measure the lattice displacement associated with the shear phonon. Even though, our findings pave the way for the investigation of phase transitions in other 2D materials, such as MoTe2, using constructive and destructive excitation of coherent shear phonons under electronic excitation.
@@
In conclusion, we demonstrate coherent control of coherent shear phonons in the time-domain using pairs of femtosecond laser pulses to examine the role of interlayer shear phonons on the lattice symmetry switching in solids. Combining double-pulse excitation and time-resolved SHG techniques in WTe2, we reveal that the Td-to-1T′ symmetry switch is independent of the coherent shear phonon motion. Furthermore, on the basis of spectroscopic ellipsometry measurements, we establish that electronic excitation, particularly interband transitions, plays a central role in the phase transition, reinforcing its classification as incoherent PIPT. These results provide a framework for understanding the role of shear phonons in structural changes in 2D materials through coherent control of interlayer phonons.

### v2

# Role of interlayer shear phonons on lattice symmetry switching in the transition metal dichalcogenide WTe2

@@ Abstract
Coherent phonon control using ultrashort pulse trains is the key to realizing structural phase transitions in solids by non-thermal pathways. By combining double-pulse excitation and time-resolved second harmonic generation techniques under high-density electronic excitation in a 2D layered material, WTe2, we demonstrate that the lattice symmetry switching from the Weyl semimetallic to the semimetallic 1 phases is independent of the amplitude of the coherent interlayer shear phonons after the arrival of the second pump pulse. This finding provides new insights into the mechanisms for symmetry switching that electronic excitation-driven shear sliding plays a dominant role.
@@
Recently, a feasible system for optical control of coherent PIPT involves ultrafast lattice symmetry switching between the semimetallic 1 (centrosymmetric) and Weyl semimetallic (non-centrosymmetric) phases of layered transition-metal dichalcogenides (TMDs) such as and [21, 22, 23, 24]. Notably, driving a coherent interlayer shear phonon displacement provides a means to control the lattice symmetry and allow periodic alternation of adjacent structural phases at sub-THz frequencies. To realize coherent PIPT for optical control, it is essential to elucidate the dynamics of incoherent and coherent PIPT in TMDs, focusing on the effect of interlayer shear displacement. Additionally, under high-density electronic excitation conditions exceeding several by a single pulse, the intended phase transition may be suppressed by sample damage or saturable absorption [25, 26, 27]. These issues can arise not only in controlling lattice symmetry but also in realizing optical control of other coherent PIPTs, such as spin states [28], and must be addressed.
In this paper, we address these issues through experimental studies on lattice symmetry change under high-density electronic excitation using time-resolved second harmonic generation (TR-SHG), a sensitive probe for tracking symmetry change [21, 22, 29, 30, 31]. By combining double-pulse excitation and TR-SHG methods in a two-dimensional layered material system, WTe2, we show that the symmetric switch from the broken-inversion Weyl semimetal phase to inversion-symmetric semimetal 1 phase is independent of the amplitude of the coherent shear phonons, which has been believed to play a central role [21].
The sample used in this study was a bulk single crystal of (001) surface -WTe2 (from HQ Graphene) with a thickness of 100 m. The sample surface was cleaved with an adhesive tape prior to measurements. SHG arises from a non-zero second-order susceptibility in materials without inversion symmetry [32, 33, 34]. In WTe2, the phase has no inversion symmetry, whereas the 1 phase does. Consequently, the phase exhibits SHG, while the 1 does not [21]. Note that when thin-layered samples are used, the contribution from the surface and/or interfaces can be prominent [35]. However, the surface contribution would appear as a background with little variation with respect to the lattice symmetry switching, since the SHG responses are allowed in both the surface and bulk regions [22, 32].
@@
Before TR-SHG measurements, static SHG measurements were performed to clarify the SHG response from -WTe2. Figure 1(b) shows the spectrum observed for the incoming probe photon (1230 nm) and the resulting SHG signal (615 nm) emitted from the sample. Figure 1(c) shows the dependence of the polarization angle on the 1230 nm light pulses measured without pump light to investigate the response of SHG to crystal orientation. SHG intensity reaches its maximum for the -axis polarization and minimum for the -axis polarization. For TR-SHG measurements, the polarization was set to the -axis to maximize the SHG intensity. Here, the intense probe (3.2 ) was used to obtain a sufficient signal-to-noise ratio due to the weak second harmonic effect. It is noted that the intense probe does not influence PIPT because the fluence was selected within the region where the intensity of SHG, , was proportional to the square of the intensity of the incident probe , i.e., up to 3.5 mJ/cm2 (see details in the Supplemental Material [36]), as expected for a second-order nonlinear process. If the probe contributes to the PIPT, the SHG should decrease following the lattice symmetry changes [37, 22]. However, this is not the case here. Furthermore, according to the absorption coefficient obtained from our spectroscopic ellipsometry measurement, the effective photogenerated carrier density can be estimated to be 4.3 times larger for the pump than for the probe at the same fluence level (see the Supplemental Material [36]). This additional measurement further supports the conclusion that the intense probe does not influence PIPT.
@@
To explore the contribution of electronic excitation and shear phonon oscillation in structural change, we performed coherent control of the interlayer shear phonon. Figure 2 shows the time-domain signal of the SHG intensity change () obtained under double-pulse excitation together with the case of single-pulse excitation at 1.5 mJ/cm2 (see the inset). As seen in the inset, the SHG intensity decreases just after photoexcitation (2 ps) and then exhibits a long-lived coherent oscillation for tens of picoseconds. The frequency of the long-lived coherent oscillation was found to be 0.24 THz, matching the interlayer shear phonon assigned as the mode [21, 38, 29]. The time interval between the first and second pulses is 1.5 (= 6.25 ps) and 2 (= 8.33 ps) of the interlayer shear phonon at the identical total pump fluence of 4.6 mJ/cm2, where = 4.17 ps is the period of the interlayer shear phonon. Note that the oscillation amplitude of the higher frequency optical mode (2.4 THz) is estimated to be 10 times smaller than that of the shear mode [39]. Therefore, the modulation of SHG intensity via the higher frequency optical mode would be too weak to be observed.
@@
The decrease in SHG intensity indicates that the lattice symmetry changes from toward 1 phases [21]. When the value of is an integral multiple of the phonon period, the oscillation amplitude of the coherent phonon generated by the first pulse is enhanced by the second pulse [11]. In contrast, when it is a half-integer multiple of the phonon period, it is canceled by the second pulse. As expected, the shear phonon was canceled when =1.5 and enhanced when =2. From Fig. 2, regardless of constructive or destructive excitation, no significant difference was observed in the decrease of the non-oscillatory component of after the arrival of the second excitation pulse (10 ps or later). Therefore, in the symmetry change from the toward 1 phases under the near-infrared (NIR) photon pump (800 nm), the effect of electronic excitation is a more dominant factor than the displacement of the shear phonon. In other words, the excitation of the coherent shear phonon can be independent of the formation of the metastable phase after the arrival of the second pump pulse (10 ps). Although decoupling of the shear phonon from the symmetry change from the toward 1 phases was suggested by suppressing the shear phonon with lattice defects [38], our results indicate more direct evidence by using coherent control of the phonon amplitude.
@@
To gain more insight into the effects of coherent control of the phonon amplitude on the possible lattice symmetry change from the toward 1 phases, we present time-domain SHG signals for both constructive ( = 8.33 ps) and destructive ( = 6.25 ps) excitation at the total pump fluence from 1.6 (1.8) mJ/cm2 to 10.6 mJ/cm2 in Fig. 3. The amplitude of the shear phonon is enhanced by constructive excitation, as shown in Fig. 3(a), whereas it is canceled by destructive excitation as shown in Fig. 3(b). In both cases, the SHG intensity decreases as the total pump fluence increases.
@@
Figure 4 shows the pump fluence dependence of the non-oscillatory obtained from the fitting in Fig. 3. As seen in Fig. 4, a higher pump fluence leads to a more significant decrease in the SHG intensity for both single- and double-pulse excitation. Moreover, even in the fluence dependence, the SHG intensity is significantly lower for double-pulse excitation than for single-pulse excitation. This suggests that saturable absorption was suppressed by multiple-pulse excitation [40, 25, 15] and/or that the structure is more likely to change if a second pump pulse is applied after the first has already modified the structure [20]. Although there has been a Landau theory to account for the temperature dependence of the SHG intensity [41], to our knowledge, an appropriate model for the fluence dependence will not currently be available under double-pulse excitation. Thus, the fluence dependence can be adapted to two linear functions with different slopes, implying a change of the potential energy surface [42, 1], e.g., a shift of the potential energy minimum toward the 1 phase, under PIPT. In this case, the critical fluence will be higher for double-pulse excitation (6 mJ/cm2) than for single-pulse excitation (4 mJ/cm2).
@@
To evaluate the impact of thermal effects during the phase transition from the Weyl semimetal to semimetal 1 phases, we have calculated the electron and lattice temperatures using the two-temperature model (TTM) [43] under double-pulse excitation as well as incorporating possible cumulative effects [44]. TTM, which applies to zero-gap metallic systems, can be used to examine the effect of elevated lattice temperatures. The TTM results for the two typical pump fluences in our experiment (7.0 mJ/cm2 and 9.2 mJ/cm2) are shown in the Supplemental Material [36]. Although the TTM analysis can only be used as a crude estimate, after the arrival of the second pulse, the lattice temperature () exceeds the transition temperature ( = 565 K [45] or 613 K [46]) at the time delay of 10 ps for 9.2 mJ/cm2. On the other hand, is below for 7.0 mJ/cm2. Interestingly, this critical fluence (7.0 mJ/cm2) is close to that observed for the depletion of the SHG intensity in Fig. 4, implying that the flatter part for 6 mJ/cm2 would need to be further addressed focusing on, in particular, early time dynamics (first few picoseconds) where the shear phonon displacements facilitate the symmetry switch. Note that the actual temperature increase estimated may slightly differ from the TTM results because the TTM maintains homogeneous absorption, no multi-photon absorption, both of which could occur under high-density photoexcitation, this is the case here.
Moreover, if SHG suppression is governed by cumulative heating, a single pulse with the same or higher fluence should yield similar suppression, but this is not the case here (Fig. 4). Thus, the cumulative heating effect would play just a minor role. During the first few picoseconds, the lattice absorbs energy, and the potential landscape is already distorted before the second (destructive) pulse arrives. The idea that destructive interference stops the shear mode tests if the larger amplitude of the shear displacement further facilitate the symmetry switch after the arrival of the second pump pulse (10 ps). The same suppressed level of the SHG signal observed in Figs. 2, 3, and 4 for constructive and destructive excitations suggest that as the prior trajectory, just after excitation by the first pump pulse, the phase already have moved toward the 1 phase. Thus, it will be required to test for the coherent control even for early time window (within a few picoseconds), although in this case we need to consider higher frequency phonons, e.g., 2.4 THz or 3.4 THz [39].
@@
We argue that the lattice symmetry switch from the to the 1 phase is driven either by interband or intraband electronic excitation. The former promotes electrons from bonding to anti-bonding states, increasing the charge carrier density . In contrast, the latter intraband transition results in an increase in the electron temperature (or electron acceleration) without changing and can preferably be achieved by irradiating far-infrared or THz pulses [21]. Both excitation processes can promote carrier (hole) doping and induce shear sliding of atomic layers toward the 1 phase [9]. Under NIR (800 nm) conditions, the photoexcited carriers in WTe2 are generated by interband transitions along and directions in the electronic band structure [47]. In fact, we measured the imaginary part () of the complex refractive index of WTe2 by spectroscopic ellipsometry, and found that the 800 nm ( 1.55 eV) pump light used in the present study was resonant with 1.56 eV (see details in the Supplemental Material [36]), which can generate the photoexcited carrier via the interband transition. Thus, the increase in charge carrier density, e.g., 1.26 1021/cm3 (at 1.6 mJ/cm2) due to interband transitions plays a dominant role in the symmetry switch.
@@
This study demonstrates that coherent control of the shear phonon amplitude can suppress its effect, isolating the impact of electronic excitation. We find that inversion-symmetry changes in the crystal structure mainly result from electronic excitation, i.e., incoherent PIPT [1]. In classical (incoherent) PIPT, a new phase emerges due to energy relaxation of electronic excited states, which refers to electron-phonon scattering process (phonon emission). The time constant for this incoherent emission will be on order of the phonon period, i.e., 4 ps for the shear mode, which is longer than the initial suppression of SHG signal (1-2 ps). Thus, the incoherent PIPT will be promoted by incoherent emission of other high-frequency phonons. Although previous attempts have used coherent shear phonons to induce structural transitions, we successfully decouple shear phonon motion from the -to-1 phase switch in WTe2 through destructive excitation after the arrival of the second pump pulse (10 ps). Since the TR-SHG technique is not a method that directly measures lattice displacement, in particular, atomic motions during this early time window (a few picoseconds), it will be important to examine the present results using direct methods such as ultrafast electron diffraction [17, 21] and time-resolved x-ray diffraction [48, 49], which can measure the lattice displacement associated with the shear phonon. Even though, our findings pave the way for the investigation of phase transitions in other 2D materials, such as MoTe2, using constructive and destructive excitation of coherent shear phonons under electronic excitation.
@@
In conclusion, we demonstrate coherent control of coherent shear phonons in the time-domain using pairs of femtosecond laser pulses to examine the role of interlayer shear phonons on the lattice symmetry switching in solids. Combining double-pulse excitation and time-resolved SHG techniques in WTe2, we reveal that the -to-1 symmetry switch is independent of the coherent shear phonon motion after the arrival of the second pump pulse. Furthermore, on the basis of spectroscopic ellipsometry measurements, we establish that electronic excitation, particularly interband transitions, plays a central role in the phase transition, reinforcing its classification as incoherent PIPT. These results provide a framework for understanding the role of shear phonons in structural changes in 2D materials through coherent control of interlayer phonons.
@@ Acknowledgement
Note added. – After this paper was submitted, we became aware of related results by Horstmann et al. [50] on precise coherent phonon control avoiding amplitude saturation.